4. Suppose your retirement goal is to accumulate $400 in 2.5 years. How much do you need to deposit in the bank today to meet your goal if the bank offers an interest rate 5.65%? a. $321 b. $336 c. $482 d. $293 e. $349
Answer #1

Required Future Value = $ 400, Tenure of Investment = 2.5 years, Interes Rate = 5.65%

Let the required deposit today be $ K

Therefore, K x (1.0565)^(2.5) = 400

K = 400 / (1.0565)^(2.5) = $ 348.647 ~ $ 349

Hence, the correct option is (e)